Fermi National Laboratory
 

Accelerator Update for March 21, 2007 - March 23, 2007

 

Wednesday March 21

The day shift began with Tevatron (TeV) in shot setup, with Operations monitoring the Recycler stash (315.1E10), with the Antiproton Source (Pbar) stacking (21.4E10), with the Booster, Linac, Main Injector (MI), and Recycler operating normally, with MiniBooNE off, with NuMI taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) waiting for the colliding beams.

Operations established store 5293 at 10:55 AM with an initial luminosity of 244.27E30.

A MI expert began a barrier bucket study at 11:45 AM.

A Pbar expert began a 30-minute target beam sweeper study at 12:44 PM.

At 1:27 PM, a Linac RF station (LRF3) was dropping its gradient. An expert investigated, reported that the driver was sparking, and fixed the problem.

At 1:34 PM, a Linac Klystron RF station (KRF5) suffered from a gun spark for the first of three times in 45 minutes. It reset.

A MI expert began a collimator study at 3:30 PM.

At 3:33 PM, all of the Linac RF station tripped off due to a safety system trip. Experts investigated. Operations loaded an Energy Conservation time line into the Time Line Generator (TLG) while the investigation continued.

Operations reported at 4 PM that the Recycler had stashed antiprotons twice during the day shift.

Linac resumed sending beam at 4:17 PM. Experts reported that a cable that was part of the low energy magenta light alert safety system in an unpowered cabinet had been unplugged. The cable was identified and reconnected.

At 4:23 PM, a Pbar Debuncher RF station (DRF1) tripped off when stacking resumed. The station had to be brought back on slowly. Stacking resumed at 4:45 PM.

The MI collimator studies ended at 5 PM.

The MI barrier bucket study ended at 6:30 PM.

At 7:30 PM, Operations reported that the Pbar stacking rate had dropped. An expert investigated and found that the Debuncher transverse cooling was set with a lower than normal power output. Stacking returned to normal at 9:02 PM.

At 9:46 PM, Operations found that the Tevatron Electron Lens (TEL1) was off. Operators contacted an expert and then turned it on and ran it up slowly.

At 9:54 PM, Operations reported that the entire beam during one MI acceleration cycle was lost; this was a slip stacked NuMI event.

At 11:42 PM, an Accumulator power supply (A:V202) alarmed due to regulation problems. Operations began a stash of antiprotons just in case this was a real problem; it wasn’t.

Operations reported at midnight that the Recycler had stashed twice during the evening shift.

Thursday March 22

The midnight shift began with Operations monitoring store 5293 (67.4E30) and the stash (225.8E10), with Pbar stacking (19.8E10), and with NuMI taking beam.

At 4:16 AM, the MI beam permit tripped off due to bad beam positions for the first of nine times in about three hours. It reset each time.

Operations prepared NuMI for an access at 7 AM. This was for scheduled safety system testing. The test lasted about 15 minutes.

Operations established slow extraction beam for a Switchyard (SY) expert to use for studies at 7:23 AM.

Operations reported at 7:59 AM that the Recycler had stashed three times during the midnight shift.

Pbar experts began a stacktail study at 8:31 AM. Stacking was halted for this study.

TeV experts changed the collision point for CDF at 10:42 AM.

Linac experts began work on LRF2 at 10:45 AM and on LRF4 at 11:22 AM.

The Pbar stacktail study ended around 12:45 PM. Normal stacking resumed.

Operations terminated store 5293 at 1:01 PM.

Linac experts completed their work on LRF2 & 4. NuMI resumed taking beam at 1:08 PM.

Cryo system technicians started work on the TeV sector E1 wet engine at 1:12 PM. This work took about 90 minutes.

Operations prepared D0 and CDF for access at 1:17 PM.

At 1:18 PM, the NuMI horn tripped off due to an interlock switch on its power supply’s door latch. It reset.

CDF came out of access at 2:15 PM.

D0 came out of access at 3:19 PM.

Operations put the TeV through a dry squeeze at 4:08 PM.

Operations reported at 4:38 PM that the Recycler had stashed once during the day shift.

Operations put the TeV into shot setup at 4:39 PM.

At 4:50 PM, the TeV suffered a sector F2 quench. Operations reported that they had been manipulating the TeV to allow an expert to look at coalescing, but had forgotten to disable a clock event that let an injection kicker move the beam away from its normal position. This delayed the shot setup.

At 5:20 PM, LRF3 tripped off. It reset.

Operations resumed shot setup at 7:29 PM.

At 9:01 PM, LRF5 tripped off on a “Cavity Pressure High” indication. Operators called in a Linac expert. This held off all downstream beam.

Operations established store 5296 at 9:11 PM with an initial luminosity of 267.52E30.

LRF5 work was completed. The expert said he had found a “flakey” dipswitch. Beam resumed at 10:40 PM.

At 10:46 PM, Operations reported that beam wasn’t making it into the Pbar Debuncher. Operators cycled the Lithium Lens power supply and this resolved the problem.

Operations reported at midnight that the Recycler had stashed once during the evening shift.

Friday March 23

The midnight shift began with Operations monitoring store 5296 (167.46E30) and the stash (75.1E10), with Pbar stacking (26.0E10), and with NuMI taking beam.

At 12:16 AM, the MI beam permit tripped off twice in four minutes due to bad beam positions. It reset.

At 12:52 AM, Operations reported receiving several alarms from the Meson MS1 and 2 areas. The Duty Electrician investigated and found that a yard breaker had tripped. He reset the breaker.

At 2:47 AM, BRF16 tripped off on an overload. Operators reset the station, but it tripped off again. Operators called in an expert and bypassed the station. At 4:14 AM, the RF expert reported that the repair would require an access.

At 5:57 AM, BRF2 tripped off on an overload. It reset.

Operations established beam at 6:21 AM to the two new MTest experiments, T950, which was the primary user, and T964, the secondary user. (T950 will measure the detector inefficiency of a particular kind of drift tube called a straw tube. T964 wants to understand the various chamber characteristics of a Gas Electron Monitor (GEM) developed at CERN. GEMs are x-ray sensitive beam monitors. The International Linear Collider intends to use them as gap monitors.)

At 6:22 AM, the MI beam permit tripped off for the first of four times in about 90 minutes. It reset each time.

At 7:14 AM, Operations reported that a Linac quadrupole’s output (L:QTM2C) had dropped to zero twice. Operators investigated and found it was a power supply thermal overload trip. Operators provided some extra cooling.

Operations reported at 7:59 AM that the Recycler had stashed three times during the midnight shift.

The Plans for Friday and the Weekend

The Run Coordinator’s plans for today are to allow a two-hour Booster access to repair three RF stations. During this access techs will clean the MI RF strainers and Pbar experts will conduct parasitic studies. Other jobs may be allowed. The plans for the weekend are to stack, stash, and store. Some studies during the weekend may be allowed.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/23/2007    email Fermilab